OpenStudy (anonymous):

Which number is an irrational number? A. The Square root of 121 B. The Square root of 150 C. The Square root of 400 D. The Square root of 1,000,000 Pls help I'm really bad at math

OpenStudy (welshfella):

3 of those numbers are perfect squares eg 121 = 11*11 find the other 2 then the other one will be the irrational number

OpenStudy (anonymous):

I don't understand square roots that much. Like I said I'm pretty bad at math. Can you explain to me?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

ok take 100 the square root is 10 because 10*10 = 100 the square root of 64 is 8 because 8*8 = 64 400 is likely to have a rational square root - what do you think it is ?

OpenStudy (welshfella):

400 ends in 0 so its square root would end in 0

OpenStudy (anonymous):

Is it 20?

OpenStudy (welshfella):

right

OpenStudy (welshfella):

now 1,000,000 is another candidate for being a perfect square the square root is 1 and how many zeros?

OpenStudy (anonymous):

3?

OpenStudy (welshfella):

right 1000*1000 = 1000000

OpenStudy (anonymous):

So the answer would be B?

OpenStudy (welshfella):

that leaves 150 which has an irrational square root

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Thank you so much can you help me with one more?

OpenStudy (welshfella):

if you factor 150 it comes to 2 * 3*5*5 = 6*25 now square root of 25 is 5 so we can write square root of 150 as 5*sqrt6 now sqrt6 is irrational ( its a decimal that never ends) so sqrt 150 is irrational

OpenStudy (welshfella):

sorry gotta go right now

OpenStudy (anonymous):

Oh Ok. Thanks for helping me though I really appreciate it! :)

OpenStudy (welshfella):

yw